# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt (eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version

### v1

@@ 4.1 Datasets and Evaluation Metrics
Table 1: The comparison of MaskSAM with SOTA methods on the AMOS testing dataset evaluated by Dice Score. To fair comparison, all results are based on 5-fold cross-validation without any ensembles. The best results are indicated as in bold.
4.2 Losses and Matching
Specifically, the desired final output $z=\{(p_{i},m_{i})\}_{i=1}^{N}$ contains $N$ pairs of binary masks $\{m_{i}^{\text{final}}|m_{i}^{\text{final}}\in[0,1]^{H\times W}\}_{i=1}^{N}$ with classes of probability distribution $p_{i}\in\Delta^{K+1}$, which contains $K$ category labels with an auxiliary ”no object“ label ($\varnothing$). Meanwhile, our model produces $N$ pairs of auxiliary boxes and masks, $z_{\text{aux}}=\{(b_{i}^{\text{aux}},m_{i}^{\text{aux}})|b_{i}^{\text{aux}}\in[0,1]^{4},m_{i}^{\text{aux}}\in\{0,1\}^{H\times W}\}^{N}_{i=1}$. Additionally, the set of $N^{gt}$ ground truth segments $z^{gt}=\{(c_{i}^{gt},b_{i}^{gt},m_{i}^{gt})|c_{i}^{gt}\in\{1,...,K\},b_{i}^{gt}\in[0,1]^{4},m_{i}^{gt}\in\{0,1\}^{H\times W}\}^{N^{gt}}_{i=1}$ is required. Since we set $N\leq N_{gt}$ and pad the set of ground truth labels with “no object token” $\varnothing$ to allow one-to-one matching. To train the model parameters, given a matching $\sigma$, the main loss $\mathcal{L}_{\text{mask-box-cls}}$ is expressed as follows,
$\mathcal{L}_{\text{mask-box-cls}}=$ $\sum_{j=1}^{N}[-\log p_{\sigma}(j)(c^{gt}_{j})+\mathds{1}_{c^{gt}_{j}\neq\varnothing}\mathcal{L}_{mask}^{\text{aux}}(m^{\text{aux}}_{\sigma}(j),m_{j}^{gt})$
$+\mathds{1}_{c^{gt}_{j}\neq\varnothing}\mathcal{L}_{box}(b_{\sigma}(j),b_{j}^{gt})+\mathds{1}_{c^{gt}_{j}\neq\varnothing}\mathcal{L}_{mask}^{\text{final}}(m^{\text{final}}_{\sigma}(j),m_{j}^{gt})].$ (1)
Method Aotra $\uparrow$ Gallbladder $\uparrow$ Kidnery(L) $\uparrow$ Kidnery(R) $\uparrow$ Liver $\uparrow$ Pancreas $\uparrow$ Spleen $\uparrow$ Stomach $\uparrow$ DSC $\uparrow$
VNet [33] 75.34 51.87 77.10 80.75 87.84 40.04 80.56 56.98 68.81
U-Net [41] 83.17 58.74 80.40 73.36 93.13 45.43 83.90 66.59 74.99
AttnUNet [35] 89.55 68.88 77.98 71.11 93.57 58.04 87.30 75.75 77.77
@@ 4.2 Losses and Matching
SAMed [48] 87.77 69.11 80.45 79.95 94.80 72.17 88.72 82.06 81.88
SAMed_s [48] 83.62 57.11 79.63 78.92 93.98 65.66 85.81 77.49 77.78
SAM3D [5] 89.57 49.81 86.31 85.64 95.42 69.32 84.29 76.11 79.56
MaskSAM (Ours) 91.75 72.20 87.32 88.15 97.21 79.62 92.47 89.11 87.23
Table 2: The comparison of MaskSAM with SOTA methods on Synapse dataset (DSC in %). The best results are indicated as in bold.
@@ 4.3 Comparison with State-of-the-Art Methods
Results on AMOS 2022 Dataset. We present the quantitative outcomes of our experiments on the AMOS 2022 Dataset Table 1, comparing our proposed MaskSAM against the segmentation methods which are widely used and well-recognized in the community, including convolution-based methods (nnUNet [24]), transformer-based methods (UNETR [17], SwinUNETR [16], and nnFormer [49]). To fair comparison, all methods run 5-fold cross-validation without any ensembles. We observe that our MaskSAM outperforms all existing methods on most organs, achieving a new state-of-the-art performance in DSC. Specifically, it surpasses nnUNet and SwinUNETR by 2.7% and 2.5% in DSC, respectively. In the extremely hard AMOS 2022 dataset, our MaskSAM achieves state-of-the-art performance, which confirms the efficacy of our method.
Method RV $\uparrow$ Myo $\uparrow$ LV $\uparrow$ DSC $\uparrow$
@@
SAM3D [5] 89.44 87.12 94.67 90.41
MaskSAM (Ours) 92.30 91.37 96.49 93.39
Table 3: The comparison of MaskSAM with SOTA methods on ACDC dataset (DSC in %). The best results are indicated as in bold.
Results on ACDC Dataset. In Table 3, we provide the quantitative experimental results on ACDC dataset. Specifically, we compare the proposed MaskSAM with several leading SAM-based method(i.e. SAM3D [5]), convolution-based methods (i.e., R50-U-Net [41] and nnUNet [24]) and transformer-based methods (i.e., TransUNet [9], SwinUNet [6], and LeViT-UNet-384s [46]). The results show that the proposed MaskSAM outperforms various state-of-the-art approaches, surpassing nnFormer by $1.3\%$, $2.3\%$, $1.8\%$, and $0.8\%$ in DSC, RV dice, Myo dice, and LV dice. Meanwhile, our method outperforms the SAM-based methods, SAM3D, by 3%, demonstrating the effectiveness of our method. In Fig. 5, we provide qualitative results compared to several state-of-the-art methods. Qualitative results depicted in Fig. 5 further illustrate the model’s accuracy across all labels, achieving near-perfect predictions in the challenging, highly saturated dataset. These outcomes affirm the efficacy of our method, as our proposed modules effectively address the limitations of SAM when adapting to medical image segmentation.
Figure 5: Qualitative comparison on the ACDC medical image segmentation dataset. MaskSAM is the most precise for each class and has fewer segmentation outliers
Results on Synapse Dataset. We present the quantitative results of our experiments in the Synapse dataset in Table 2, comparing our proposed MaskSAM against several leading SAM-based method(i.e., SAMed [48] and SAM3D [5]), convolution-based methods (VNet [41] and nnUNet [24]), transformer-based methods (TransUNet [9], SwinUNet [6], and nnFormer [49]). We observe that MaskSAM outperforms all existing methods, achieving a new state-of-the-art performance. Specifically, it surpasses nnFormer by 0.7% in the DSC for the highly saturated dataset. Meanwhile, our method outperforms the SAM-based methods, SAMed and SAM3D, by more than 5%, demonstrating the effectiveness of our method. Notably, our model excels in predicting the large-size ‘Liver,’ ‘Spleen,’ and ‘Stomach’ labels, attributed to our innovative DConvAdapter and DMLPAdapter. These adapters enable the learning of more intricate 3D spatial information and adapt 2D SAM to medical image segmentation. In Fig. 6, we illustrate qualitative results compared to representative methods. These results also demonstrate that our MaskSAM can predict more accurately the ‘Liver’, ‘Spleen’, and ‘Stomach’ labels. In conclusion, the effectiveness of our method is robustly demonstrated by both quantitative and qualitative results.
Figure 6: Qualitative comparison on the Synapse medical image segmentation dataset. MaskSAM is the most precise for each class and has fewer segmentation outliers.
@@ 4.4 Ablation Study
Baseline Models. The proposed MaskSAM has 9 baselines (i.e., B1, B2, B3, B4, B5, B6, B7, B8, B9) as shown in Table 4. All baselines contain the entire SAM structure, a prompt generator, and a learnable class token. (i) B1 adopts the prompt generator that only generates learnable binary masks and uses the binary masks as mask prompts to calculate its bounding boxes as box prompts shown in Fig. 2(a). (ii) B2 adopts the prompt generator that only generates learnable boxes as the prompts shown in Fig. 2(b). (iii) B3 adopts the prompt generator that generates learnable binary masks as mask prompts and learnable boxes as box prompts shown in Fig. 2(c). (iv) B4 adopts the prompt generator that generates learnable binary masks as mask prompts and learnable boxes. We average the bounding boxes calculated from binary masks and the learnable boxes as the final box prompts shown in Fig. 2(d). (v) B5 adds depth positional embedding blocks (DPosEmbed) in the image encoder and mask decoder based on B4. (vi) B6 modifies the vanilla adapter by inserting the invert-bottleneck depth MLPs with a skip connection after the fully-connected layers for upsampling based on B5. (vii) B7 modifies the vanilla adapter by inserting the invert-bottleneck depth MLPs with a skip connection before the fully-connected layers for downsampling based on B5. (viii) B8 replaces the vanilla adapter with our designed DMLPAdapter based on B5. (ix) B9 is our full model, named MaskSAM, illustrated in Fig. 1. B9 adopts the DMLPAdapter for prompt embeddings and the DConvAdapter for image embedding based on B8. The results of the ablation study are shown in Table 4.
Ablation analysis. The results of the ablation study are shown in Table 4. When we use our proposed MaskAvgBBoxPG to first generate auxiliary binary masks and auxiliary boxes, then average the bounding boxes calculated from auxiliary binary masks and learnable auxiliary boxes as the final box prompts, the model achieves the best results and improves by 1.9%, 2.6% and 1.3% compared to B1 with a learnable mask prompt generator, B2 with a learnable box prompt generator, and B3 with a learnable mask and a learnable box prompt generator, respectively. The result confirms the effectiveness of the proposed prompt generator. When inserting depth positional embedding (DPosEmbed) into the image encoder and mask decoder, the performance of B5 improves by more than 0.15% compared to B4, demonstrating the effectiveness of DPosEmbed blocks. The DMLPAdapter (B8) which we insert depth MLPs with a skip connection in the middle of the vanilla adapter achieves the best performance and improves by 0.2% and 0.1% compared to B6 in which we insert depth MLPs with a skip connection after fully-connected layers for upsampling and B7 in which depth MLPs are inserted with a skip connection after fully-connected layers for downsampling, respectively. The result confirms the effectiveness of the proposed DMLPAdapter. B9 is our full model, MaskSAM, utilizing the DMLPAdapter for prompt embeddings and the DConvAdapter that we replace the depth MLP with a 3D depth-wise convolution layer from the DMLPAdapter for image embedding based on B8 as shown in Fig. 1. Compared to B8, our model brings about 0.3% improvements. Therefore, the results demonstrate the effectiveness of our proposed MaskSAM.
@@
Table 4: The ablation studies of the proposed method on the ACDC dataset. The {}PG means a prompt generator with different prompts. The vAdapter means the vanilla adapter. The D-MLP means the MLP layers processing on the depth dimension. The DPosEmbed means the depth positional embedding.

### v2

@@ 2.1 Overview of the Proposed MaskSAM
Table 1: The comparison of MaskSAM with SOTA methods on the AMOS testing dataset evaluated by Dice Score. To fair comparison, all results are based on 5-fold cross-validation without any ensembles. “Semantic labels” indicate the model’s ability for semantic labeling, while “Prompt” specifies the source of the prompt. The best results are indicated as in bold.
2.2 Proposed Dataset Mapping
SAM generates a single binary mask without any semantic label for each prompt, while a typical ground truth of medical images comprises multiple classes. Each cropped patch image as input may contain varied classes. The challenge arises from varying the lengths of binary masks within the ground truth. To accommodate this diversity and inspired by DETR [7] and MaskFormer [9], our model generates a sufficiently large number of binary masks, with each mask dedicated to predicting a single class. And then bipartite matching is utilized between the set of predictions and ground truth segments. Therefore, we design a dataset mapping pipeline, shown at the bottom of Figure 2, which converts a multi-class mask into a set of binary masks with semantic labels per class.
2.3 Proposed Prompt Generator
To solve the requirements of the extra proper prompts, we propose a prompt generator, shown in Figure 3(b), to generate a set of auxiliary binary masks and bounding boxes as prompts instead of manual prompts for the prompt encoder. We adopt box prompts with mask prompts as our prompt way, since the point prompts would bring instabilities that are harmful to medical segmentation tasks. We utilize ViT’s strong representation capabilities to extract multiple levels of feature maps from the image encoder as the input of our proposed prompt generator. From the beginning of the last output of the image encoder, the feature maps are connected with convolution layers, upsampled, and concatenated with the feature maps of the lower level. Following this way, we can obtain feature maps with the same image size as the ground truth. Finally, we utilize a convolutional layer to change the channel size to the fixed number of $N$ that is larger than the maximum of object-level binary masks in datasets. Meanwhile, we extract the output of the last convolutional layer at each level, use adaptive average pooling layers to change the spatial dimension to $(2,2)$ for box queries, concatenate all box queries, and connect with an MLP layer to adjust the channel to $N$. We can obtain $N$ number of learnable binary masks and learnable boxes.
There are many combinations of learnable binary masks and learnable boxes shown in Figure 5(a)-(d). Figure 5(a) shows a prompt generator only generates learnable binary masks and uses the binary masks to calculate its bounding boxes. Figure 5(b) shows a prompt generator only generates boxes as prompts. Figure 5(c) shows a prompt generator that generates learnable binary masks as mask prompts and learnable boxes as box prompts. Figure 5(d) shows a prompt generator that generates learnable binary masks as mask prompts and learnable boxes. We average the bounding boxes calculated from binary masks and the learnable boxes as the final box prompts. Through a series of experiments, we found that the best way is Figure 5(d) since it can involve more information and more robustness.
Figure 6: Qualitative comparison on Synapse dataset. MaskSAM is the most precise for each class and has fewer segmentation outliers.
2.4 Proposed Adapters
we adopt the lightweight adapter [20] which is a bottleneck architecture that consists of two fully connected (FC) layers and an activation layer in the middle to modify and inject into each transformer block during fine-tuning. Unlike classic 2D natural images, many medical scans are 3D volumes with extra depth dimensions such as MRI and CT. To learn extra depth information, we involve learnable layers for the extra depth dimension into adapters. In SAM, the image encoder and mask decoder contain transformer blocks that we can insert adapters into. The mask decoder contains two types of attention blocks for prompt embeddings and image embeddings, respectively. The original adapter only processes the last dimension, the channel dimension, which cannot learn the information among the tokens. However, image embeddings contain spatial information, which is important to let our model understand spatial relationships. we design a 3D depth-convolution adapter (DConvAdapter), shown in Figure 4(b), that adds a 3D depth-wise convolution layer in the middle of the original adapter with a skip connection for all attention blocks processing image embeddings in the mask decoder. For the rest of the attention blocks for prompt embeddings, we only need to involve a learnable block at depth dimension, since the prompt embeddings do not have any spatial relationship. Therefore, we design a 3D depth-MLP adapter (DMLPAdapter), shown in Figure 4(a), that adds an invert-bottleneck architecture that consists of two FC layers and an activation layer processing in the depth dimension in the middle of the original adapter with a skip connection, which can learn additional depth information. Since the image encoder only contains the attention blocks for the image embeddings, we insert the DConvAdapter for each transformer block into the image encoder. Figure 4(c) shows the way we insert adapters into vision transformers. We insert an adapter behind the multi-head attention block and parallel MLP block.
2.5 Modified Image Encoder
Figure 3(a) illustrates the redesigned image encoder. i) SAM works on natural images that have 3 channels for RGB while medical images have varied modalities as channels. There are gaps between the varied modalities of medical images and the RGB channels of natural images. Therefore, we design a sequence of convolutional layers to an invert-bottleneck architecture to learn the adaption from the varied modalities with any size to 3 channels. ii) The image encoder includes one positional embedding. To better understand the extra depth information, we can insert a learnable depth positional embedding with the original positional embedding. iii) Since we use the base ViT backbone, it contains 12 attention blocks. We insert our designed DConvAdapter blocks into each attention block. We extract the feature maps of each three attention blocks and the final output of the image encoder for the prompt generator.
2.6 Modified Mask Decoder
Figure 3(c) illustrates the redesigned mask encoder. i) we design learnable global classifier tokens, which are summed by auxiliary classifier tokens generated by the prompt generator, concatenated with sparse prompt embedding, and the original mask tokens to equip our model with the functionality to predict semantic labels for each binary mask. ii) The mask decoder also includes a positional embedding. To learn the extra depth information better, we can insert a learnable depth positional embedding with the original image positional embedding. iii) The mask encoder contains two subsequent transformers. Each transformer first applies self-attention to the prompt embedding. We insert a DMLPAdapter behind the self-attention. Then, a cross-attention block is adopted for tokens attending to image embedding. We insert a DMLPAdapter behind the cross-attention. Next, we insert a DMLPAdapter parallel to an MLP block. Finally, a cross-attention block is utilized for image embedding attending to tokens. We insert a DConvAdapter behind the cross-attention. In this way, our model can learn spatial information with extra depth information for the image embedding and depth information for the prompt embedding.
Sem. Lab. Prompts Method Aorta GB L.Kd R.Kd Liv. Panc. Spl. Stom. DSC
@@ 2.6 Modified Mask Decoder
GT SAM [26] bbox 60.05 24.90 68.87 54.22 76.91 45.36 69.20 67.93 58.43
GT MedSAM [31] bbox 70.16 22.44 79.08 64.63 76.38 52.31 73.07 79.10 64.65
GT MaskSAM (Ours) bbox 96.22 93.31 94.75 94.25 97.69 89.13 96.68 95.58 94.70
No need SAMed [44] 87.77 69.11 80.45 79.95 94.80 72.17 88.72 82.06 81.88
No need SAMed_s [44] 83.62 57.11 79.63 78.92 93.98 65.66 85.81 77.49 77.78
No need SAM3D [5] 89.57 49.81 86.31 85.64 95.42 69.32 84.29 76.11 79.56
No need MaskSAM (Ours) 91.75 72.20 87.32 88.15 97.21 79.62 92.47 89.11 87.23
Table 2: Comparisons with SOTA methods on Synapse.
2.7 Losses and Matching
Specifically, the desired final output $z=\{(p_{i},m_{i})\}_{i=1}^{N}$ contains $N$ pairs of binary masks $\{m_{i}^{\text{final}}|m_{i}^{\text{final}}\in[0,1]^{H\times W}\}_{i=1}^{N}$ with classes of probability distribution $p_{i}\in\Delta^{K+1}$, which contains $K$ category labels with an auxiliary ”no object“ label ($\varnothing$). Meanwhile, our model produces $N$ pairs of auxiliary boxes and masks, $z_{\text{aux}}=\{(b_{i}^{\text{aux}},m_{i}^{\text{aux}})|b_{i}^{\text{aux}}\in[0,1]^{4},m_{i}^{\text{aux}}\in\{0,1\}^{H\times W}\}^{N}_{i=1}$. Additionally, the set of $N^{gt}$ ground truth segments $z^{gt}=\{(c_{i}^{gt},b_{i}^{gt},m_{i}^{gt})|c_{i}^{gt}\in\{1,...,K\},b_{i}^{gt}\in[0,1]^{4},m_{i}^{gt}\in\{0,1\}^{H\times W}\}^{N^{gt}}_{i=1}$ is required. Since we set $N\geq N_{gt}$ and pad the set of ground truth labels with “no object token” $\varnothing$ to allow one-to-one matching. To train the model parameters, given a matching $\sigma$, the main loss $\mathcal{L}_{\text{mask-box-cls}}$ is expressed as follows, $\mathcal{L}_{\text{mask-box-cls}}=$
$\sum_{j=1}^{N}[-\log p_{\sigma}(j)(c^{gt}_{j})+\mathds{1}_{c^{gt}_{j}\neq\varnothing}\mathcal{L}_{mask}^{\text{aux}}(m^{\text{aux}}_{\sigma}(j),m_{j}^{gt})$
$+\mathds{1}_{c^{gt}_{j}\neq\varnothing}\mathcal{L}_{box}(b_{\sigma}(j),b_{j}^{gt})+\mathds{1}_{c^{gt}_{j}\neq\varnothing}\mathcal{L}_{mask}^{\text{final}}(m^{\text{final}}_{\sigma}(j),m_{j}^{gt})].$
3 Experiments
Figure 7: Qualitative comparison on ACDC dataset. MaskSAM is the most precise for each class and has fewer segmentation outliers
Datasets and Evaluation Metrics. We use three publicly available datasets, AMOS22 Abdominal CT Organ Segmentation [25], Synapse multiorgan segmentation [27] and Automatic Cardiac Diagnosis Challenge (ACDC) [3]. (i) AMOS22 dataset consists of 200 cases of abdominal CT scans with 16 anatomies manually annotated for abdominal multi-organ segmentation. There are 200 testing images, and we evaluate our model on the AMOS22 leaderboard. (ii) Synapse dataset consists of 30 cases of abdominal CT scans. Following the split strategies [8], we use a random split of 18 training cases and 12 cases for validation. We evaluate the model performance via the average Dice score (DSC) on 8 abdominal organs. (iii) ACDC dataset consists of 100 patients with the labels involving the right ventricle (RV), myocardium (MYO), and left ventricle (LV). We use a random split of 70 training cases, 10 validation cases, and 20 testing cases. We evaluate the model performance by the average DSC. In Tab. 1, 2, 3, “Semantic labels” indicate the model’s ability for semantic labeling, while “Prompt” specifies the source of the prompt. Since SAM and MedSAM do not predict semantic labels and need the extra prompts, we utilize GT or the inferred prediction by the pre-trained nnUNet to generate prompts and use the same predictions’ labels as the semantic labels.
@@ 3.1 Comparison with State-of-the-Art Methods
Results on AMOS 2022 Dataset. We present the quantitative outcomes of our experiments on the AMOS 2022 Dataset Tab. 1, comparing our proposed MaskSAM against the segmentation methods which are widely used and well-recognized in the community, including convolution-based methods (nnUNet [22]), transformer-based methods (UNETR [17], SwinUNETR [16], and nnFormer [46]), and SAM-based methods (SAM [26], MedSAM [31], SAMed [44], and SAM3D [5]). To fair comparison, all methods run 5-fold cross-validation without any ensembles. We observe that our MaskSAM outperforms all existing methods on most organs, achieving a new state-of-the-art performance in DSC. Meanwhile, the performance of points as the prompts are worse than boxes as the prompts. When utilizing the predictions from nnUNet for bounding box prompts, SAM and MedSAM exhibit decreases of 34% and 27%, respectively, compared to nnUNet’s accuracy of 87.8%. These reductions in accuracy indicate negative implications for the results. Specifically, MaskSAM surpasses nnUNet and SwinUNETR by 2.7% and 2.5% in DSC, respectively. MaskSAM surpasses SAMed and SAM3D by 13% and 15% in DSC, respectively. In the extremely hard AMOS 2022 dataset, our MaskSAM achieves state-of-the-art performance, which confirms the efficacy of our method.
Semantic labels Prompts Method RV $\uparrow$ Myo $\uparrow$ LV $\uparrow$ DSC $\uparrow$
@@
GT SAM [26] bbox 70.95 33.33 81.31 62.53
GT MedSAM [31] bbox 85.86 80.31 92.33 86.17
GT MaskSAM bbox 94.27 91.69 96.92 94.30
No needs SAM3D [5] 89.44 87.12 94.67 90.41
No needs MaskSAM (Ours) 92.30 91.37 96.49 93.39
Table 3: The comparison of MaskSAM with SOTA methods on ACDC dataset (DSC in %). The best are in bold.
Results on ACDC Dataset. In Tab. 3, we provide the quantitative experimental results on ACDC dataset. Specifically, we compare the proposed MaskSAM with several leading SAM-based method(i.e. SAM3D [4]), convolution-based methods (i.e., nnUNet [22]) and transformer-based methods (i.e., UNETR [17] and nnFormer [46]). The results show that the proposed MaskSAM outperforms various state-of-the-art approaches, surpassing nnFormer by $1.3\%$, $2.3\%$, $1.8\%$, and $0.8\%$ in DSC, RV dice, Myo dice, and LV dice. Meanwhile, our method outperforms the SAM-based methods, SAM (1 bbox), MedSAM (1 bbox), and SAM3D, by 31%, 7%, 6%, and 3%, demonstrating the effectiveness of our method. In Fig. 7, we provide qualitative results compared to several state-of-the-art methods. Qualitative results depicted in Fig. 7 further illustrate the model’s accuracy across all labels, achieving near-perfect predictions in the challenging, highly saturated dataset. These outcomes affirm the efficacy of our method, as our proposed modules effectively address the limitations of SAM when adapting to medical image segmentation.
Results on Synapse Dataset. We present the quantitative results of our experiments in the Synapse dataset in Tab. 2, comparing our proposed MaskSAM against several leading SAM-based method(i.e., SAMed [44] and SAM3D [4]), convolution-based methods (VNet [37] and nnUNet [22]), transformer-based methods (TransUNet [8], SwinUNet [6], and nnFormer [46]). We observe that MaskSAM outperforms all existing methods, achieving a new state-of-the-art performance. Specifically, it surpasses nnFormer by 0.7% in the DSC for the highly saturated dataset. Meanwhile, our method outperforms the SAM-based methods, SAM (1 bbox), MedSAM (1 bbox), SAMed, and SAM3D, by 29%, 23%, 6%, and 8%, demonstrating the effectiveness of our method. Notably, our model excels in predicting the large-size ‘Liver,’ ‘Spleen,’ and ‘Stomach’ labels, attributed to our innovative DConvAdapter and DMLPAdapter. These adapters enable the learning of more intricate 3D spatial information and adapt 2D SAM to medical image segmentation. In Fig. 6, we illustrate qualitative results compared to representative methods. These results also demonstrate that our MaskSAM can predict more accurately the ‘Liver’, ‘Spleen’, and ‘Stomach’ labels. In conclusion, the effectiveness of our method is robustly demonstrated by both quantitative and qualitative results.
@@ 3.2 Ablation Study
Baseline Models. The proposed MaskSAM has 9 baselines (i.e., B1, B2, B3, B4, B5, B6, B7, B8, B9) as shown in Tab. 4. All baselines contain the entire SAM structure, a prompt generator, and a learnable class token. (i) B1 adopts the prompt generator that only generates learnable binary masks and uses the binary masks as mask prompts to calculate its bounding boxes as box prompts shown in Fig. 5(a). (ii) B2 adopts the prompt generator that only generates learnable boxes as the prompts shown in Fig. 5(b). (iii) B3 adopts the prompt generator that generates learnable binary masks as mask prompts and learnable boxes as box prompts shown in Fig. 5(c). (iv) B4 adopts the prompt generator that generates learnable binary masks as mask prompts and learnable boxes. We average the bounding boxes calculated from binary masks and the learnable boxes as the final box prompts shown in Fig. 5(d). (v) B5 adds depth positional embedding blocks (DPosEmbed) in the image encoder and mask decoder based on B4. (vi) B6 modifies the vanilla adapter by inserting the invert-bottleneck depth MLPs with a skip connection after the fully-connected layers for upsampling based on B5. (vii) B7 modifies the vanilla adapter by inserting the invert-bottleneck depth MLPs with a skip connection before the fully-connected layers for downsampling based on B5. (viii) B8 replaces the vanilla adapter with our designed DMLPAdapter based on B5. (ix) B9 is our full model, named MaskSAM, illustrated in Fig. 2. B9 adopts the DMLPAdapter for prompt embeddings and the DConvAdapter for image embedding based on B8.
@@
Table 4: Ablation studies of proposed methods on ACDC. {}PG means a prompt generator with different prompts. vAdapter means vanilla adapter. D-MLP means MLP layers on depth dimension. DPosEmbed means depth positional embedding.
Ablation analysis. The results of the ablation study are shown in Tab. 4. When we use our proposed MaskAvgBBoxPG to first generate auxiliary binary masks and auxiliary boxes, then average the bounding boxes calculated from auxiliary binary masks and learnable auxiliary boxes as the final box prompts, the model achieves the best results and improves by 1.9%, 2.6% and 1.3% compared to B1 with a learnable mask prompt generator, B2 with a learnable box prompt generator, and B3 with a learnable mask and a learnable box prompt generator, respectively. The result confirms the effectiveness of the proposed prompt generator. When inserting depth positional embedding (DPosEmbed) into the image encoder and mask decoder, the performance of B5 improves by more than 0.15% compared to B4, demonstrating the effectiveness of DPosEmbed blocks. The DMLPAdapter (B8) which we insert depth MLPs with a skip connection in the middle of the vanilla adapter achieves the best performance and improves by 0.2% and 0.1% compared to B6 in which we insert depth MLPs with a skip connection after fully-connected layers for upsampling and B7 in which depth MLPs are inserted with a skip connection after fully-connected layers for downsampling, respectively. The result confirms the effectiveness of the proposed DMLPAdapter. B9 is our full model, MaskSAM, utilizing DMLPAdapter for prompt embeddings and DConvAdapter that we replace the depth MLP with a 3D depth-wise convolution layer from the DMLPAdapter for image embedding based on B8 as shown in Fig. 2. Compared to B8, our model brings about 0.3% improvements. Therefore, the results demonstrate the effectiveness of our proposed MaskSAM.
# Params (M) nnUNet MedSAM SAMed Med-SA
Tunable Params 29M 91M 19M 13M
Total Params 29M 91M 636M+19M 636M+13M
# Params (M) SAM3D 3DSAM-Adapter AutoProSAM MaskSAM (ours)
Tunable Params 2M 26M 27M 14M
Total Params 91M+2M 91M+26M 91M+27M 91M+14M
Table 5: Comparisons of tunable and total parameters.
3.3 Model Analysis
Efficiency metrics. Tab. 5 compares model efficiency by tunable and total parameters. Our method, with only 14M tunable parameters, effectively adapts the 91M ViT-base SAM model and ensures efficiency.
Box prompt visualization. Fig. 1 shows generated box prompts. For complex objects and images with a few objects, our model tends to generate multiple prompts to assist in mask prediction, as the object queries outnumber the objects in 2D images. The rest of the queries are assigned as no-objects. Additionally, the generated boxes are typically larger than the objects themselves, allowing the model to account for adjacent objects. As a result, the generated auxiliary box prompts are both precise and meaningful.